OpenStudy (anonymous):

What famous Pythagorean triple has values G, 21, and 29?

Parth (parthkohli):

A Pythagorean triplet is in form a,b,c where \(a^2 + b^2 = c^2\)

Parth (parthkohli):

So, \(g^2 + 21^2 = 29^2\) Just solve for g.

OpenStudy (anonymous):

a^2+441=841 Subtract 441 on both sides and get sqrt"400". Once you do the square root of both sides, it becomes a=20.

OpenStudy (anonymous):

LOL, Im an old guy who has forgotten his math,,,,how do I solve for g

Parth (parthkohli):

21^2 = 441 29^2 = 841. \( \color{Black}{\Rightarrow g^2 + 441 = 841 }\) \( \color{Black}{\Rightarrow g^2 = 400}\) \( \color{Black}{\Rightarrow g = \sqrt{400}}\) 400 is a perfect square I believe ^_^

OpenStudy (anonymous):

Yep, it goes down to 20.

Parth (parthkohli):

Exactly. :D

OpenStudy (anonymous):

I'm lost, I'm sorry so g=400?

Parth (parthkohli):

No..

Parth (parthkohli):

\( \color{Black}{\Rightarrow g^2 = 400 }\) If you find the square root of both sides: \( \color{Black}{\Rightarrow g = \sqrt{4000} }\)

Parth (parthkohli):

\(g = \sqrt{400}\) correction

OpenStudy (anonymous):

so g is the square root of 400

Parth (parthkohli):

Yeah, but square root of 400 is an integer :)